# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ B.1 Results after hyperparameter optimization
Table 14: Average test ROC-AUC per dataset for dataset-specific neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by "-".
@@
Dataset FT-Transformer ResNet SAINT TabNet adult 0.914869 0.913790 0.920246 0.882450 analcatdata_authorship 0.999985 1.000000 0.999974 0.999249 analcatdata_dmft 0.576947 0.584338 0.544695 0.515962 balance-scale 0.999735 0.989061 0.999266 0.979668 bank-marketing 0.938198 0.935740 0.936560 0.887319 banknote-authentication 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.820159 0.850801 - - blood-transfusion-service-center 0.745975 0.738502 0.746726 0.660675 breast-w 0.989503 0.995477 0.988470 0.986694 car 0.999751 0.994154 1.000000 1.000000 churn 0.914596 0.918713 0.915603 0.891443 climate-model-simulation-crashes 0.934671 0.918990 0.925643 0.868204 cmc 0.739402 0.737829 0.738490 0.647121 cnae-9 0.994497 0.997106 - - connect-4 0.901170 0.933333 - - credit-approval 0.935798 0.933113 0.933493 0.878500 credit-g 0.783048 0.783524 0.786402 0.696905 cylinder-bands 0.915494 0.909989 0.923391 0.837792 diabetes 0.831108 0.821798 0.827285 0.756416 dna 0.990937 0.992543 0.992473 0.991448 dresses-sales 0.620033 0.575205 0.624704 0.555993 electricity 0.963076 0.960658 0.967012 0.938656 eucalyptus 0.923933 0.916785 0.925970 0.872365 first-order-theorem-proving 0.796707 0.784636 0.802392 0.774094 GesturePhaseSegmentationProcessed 0.895166 0.914196 0.919006 0.850596 har 0.999685 0.999921 - 0.999515 ilpd 0.751488 0.747491 0.698718 0.704840 Internet-Advertisements 0.974513 0.974187 - - isolet 0.998817 0.999401 - 0.998813 jm1 0.709321 0.720444 0.719464 0.674043 jungle_chess_2pcs_raw_endgame_complete 0.999975 0.999956 0.999926 0.991981 kc1 0.783519 0.806819 0.796918 0.762807 kc2 0.832014 0.833248 0.834436 0.713458 kr-vs-kp 0.999777 0.999369 0.999789 0.998872 letter 0.999919 0.999926 0.999853 0.999606 madelon 0.747391 0.605018 - 0.630669 mfeat-factors 0.999015 0.999472 0.999385 0.998125 mfeat-fourier 0.984511 0.981725 0.980508 0.970539 mfeat-karhunen 0.998682 0.998448 0.999078 0.996960 mfeat-morphological 0.970198 0.968651 0.967681 0.955818 mfeat-pixel 0.997451 0.998690 0.999217 0.998200 mfeat-zernike 0.983479 0.984488 0.981874 0.968629 MiceProtein 0.999973 0.999973 1.000000 0.999344 nomao 0.990908 0.993048 - - numerai28.6 0.530315 0.528012 0.525822 - optdigits 0.999616 0.999927 0.999841 0.998871 ozone-level-8hr 0.919484 0.925416 0.919315 0.864067 pc1 0.917591 0.889458 0.870543 0.804412 pc3 0.828743 0.829637 0.827322 0.788151 pc4 0.934944 0.944447 0.934528 0.920943 pendigits 0.999703 0.999638 0.999782 0.999753 PhishingWebsites 0.996760 0.996975 0.996746 0.996196 phoneme 0.965071 0.963591 0.960382 0.956279 qsar-biodeg 0.919584 0.932220 0.930632 0.902748 satimage 0.993516 0.991995 0.992630 0.987482 segment 0.994124 0.993581 0.994831 0.992317 semeion 0.995548 0.997689 0.997630 0.994019 sick 0.997937 0.968841 0.998281 0.981838 spambase 0.985969 0.987683 0.986263 0.980804 splice 0.992276 0.993514 0.995073 0.990441 steel-plates-fault 0.959182 0.949067 0.955379 0.947456 texture 0.999983 0.999999 0.999976 0.999763 tic-tac-toe 0.996152 0.999462 0.999725 0.993030 vehicle 0.963362 0.967212 0.955127 0.943787 vowel 0.999713 0.999813 0.999875 0.999686 wall-robot-navigation 0.999900 0.999042 0.999844 0.997585 wdbc 0.993967 0.995409 0.995546 0.986656 wilt 0.993047 0.990726 0.993139 0.991289
Table 15 shows the raw results after HPO for MLP with PLR embeddings and RealMLP.
Table 15: Average test ROC-AUC per dataset for MLP and RealMLP after hyperparameter optimization across CV folds. Missing datasets are represented by "-".
Dataset MLP RealMLP adult 0.928689 0.923327 analcatdata_authorship 0.999770 1.000000 analcatdata_dmft 0.574532 0.574396 balance-scale 0.998659 1.000000 bank-marketing 0.937054 0.937031 banknote-authentication 1.000000 1.000000 Bioresponse 0.825631 0.859065 blood-transfusion-service-center 0.770627 0.746350 breast-w 0.992380 0.992882 car 0.999992 1.000000 churn 0.922938 0.913533 climate-model-simulation-crashes 0.948857 0.962163 cmc 0.744580 0.735472 cnae-9 0.996716 0.997569 connect-4 0.927373 0.928258 credit-approval 0.938866 0.917352 credit-g 0.788476 0.779381 cylinder-bands 0.886405 0.910680 diabetes 0.837342 0.837507 dna 0.992220 0.994111 dresses-sales 0.635468 0.537849 electricity 0.969201 0.961467 eucalyptus 0.921873 0.915693 first-order-theorem-proving 0.798812 0.795637 GesturePhaseSegmentationProcessed 0.911434 0.901441 har 0.999783 0.999959 ilpd 0.671938 0.729412 Internet-Advertisements - 0.973810 isolet 0.998295 0.999635 jm1 0.715620 0.713988 jungle_chess_2pcs_raw_endgame_complete 0.999965 0.999774 kc1 0.805465 0.796117 kc2 0.829426 0.845768 kr-vs-kp 0.999686 0.999704 letter 0.999894 0.999914 madelon 0.883991 0.930302 mfeat-factors 0.998875 0.999625 mfeat-fourier 0.984929 0.985483 mfeat-karhunen 0.998849 0.999019 mfeat-morphological 0.967719 0.969994 mfeat-pixel 0.998674 0.999492 mfeat-zernike 0.984610 0.982993 MiceProtein 0.999973 0.999971 nomao 0.986577 0.989803 numerai28.6 0.525920 0.529534 optdigits 0.999794 0.999968 ozone-level-8hr 0.927900 0.923252 pc1 0.832532 0.844517 pc3 0.842511 0.814590 pc4 0.945813 0.939257 pendigits 0.999705 0.999659 PhishingWebsites 0.996991 0.997208 phoneme 0.967617 0.966456 qsar-biodeg 0.924951 0.929226 satimage 0.992308 0.993034 segment 0.995046 0.994075 semeion 0.997350 0.998976 sick 0.997048 0.998661 spambase 0.988185 0.987799 splice 0.994053 0.994420 steel-plates-fault 0.964693 0.959639 texture 0.999991 0.999999 tic-tac-toe 1.000000 0.999711 vehicle 0.961813 0.965844 vowel 0.999638 0.999955 wall-robot-navigation 0.999689 0.998720 wdbc 0.996065 0.996038 wilt 0.997690 0.993197
@@
Table 16: Average test ROC-AUC per dataset for meta-learned neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by "-".
@@
Dataset CARTE TPBerta TabPFN XTab adult 0.902677 - - - analcatdata_authorship 0.999181 - 1.000000 0.999991 analcatdata_dmft 0.586376 - 0.586630 0.556971 balance-scale 0.999413 - 0.997656 0.997420 bank-marketing 0.924664 - - - banknote-authentication 1.000000 0.994512 - 1.000000 Bioresponse - - - - blood-transfusion-service-center 0.739571 0.633041 0.752586 - breast-w 0.987912 0.986514 0.994131 0.989666 car 0.997126 - - - churn 0.923626 - - - climate-model-simulation-crashes 0.938531 - 0.968010 0.944367 cmc 0.738379 - - - cnae-9 0.990151 - - - connect-4 - - - - credit-approval 0.909279 0.901989 0.932397 0.939620 credit-g 0.769619 - 0.768476 - cylinder-bands 0.848539 0.820399 0.886616 0.881396 diabetes 0.823615 0.778356 0.836120 0.815847 dna 0.986120 - - 0.992479 dresses-sales 0.589655 0.534893 0.538916 0.613136 electricity 0.909407 - - 0.966899 eucalyptus 0.905245 - 0.928493 0.918317 first-order-theorem-proving 0.764092 - - 0.798803 GesturePhaseSegmentationProcessed 0.798024 - - 0.886960 har - - - - ilpd 0.704712 0.586083 0.757892 0.726413 Internet-Advertisements - - - - isolet - - - - jm1 0.728512 - - 0.727984 jungle_chess_2pcs_raw_endgame_complete 0.973383 - - 0.999950 kc1 0.797680 - - 0.803082 kc2 0.842828 - 0.850065 0.835476 kr-vs-kp 0.999685 0.855273 - 0.999616 letter 0.999440 - - 0.999859 madelon 0.836760 - - 0.845746 mfeat-factors 0.996064 - - 0.998443 mfeat-fourier 0.976986 - - 0.982539 mfeat-karhunen 0.994814 - - 0.998582 mfeat-morphological 0.967325 - - 0.967136 mfeat-pixel 0.996175 - - 0.998642 mfeat-zernike 0.978119 - - 0.980183 MiceProtein 0.999582 - - 1.000000 nomao - - - 0.992727 numerai28.6 0.514361 - - 0.528062 optdigits 0.999112 - - 0.999712 ozone-level-8hr 0.890063 - - 0.915744 pc1 0.835444 - - 0.855741 pc3 0.831574 0.625642 - 0.823532 pc4 0.937337 0.744304 - 0.938455 pendigits 0.999468 - - 0.999751 PhishingWebsites 0.994582 - - 0.996896 phoneme 0.948702 0.796404 - 0.961749 qsar-biodeg 0.921153 0.833852 - 0.926795 satimage 0.988038 - - 0.992918 segment 0.993491 - - 0.994697 semeion 0.993378 - - 0.997064 sick 0.995762 - - 0.998232 spambase 0.983228 - - 0.986044 splice 0.987950 - - 0.992444 steel-plates-fault 0.943636 - - 0.957088 texture 0.999541 - - 0.999962 tic-tac-toe 0.984361 0.993803 0.996086 1.000000 vehicle 0.941691 - 0.970556 0.955838 vowel 0.998092 - - 0.999630 wall-robot-navigation 0.999505 - - 0.999846 wdbc 0.990612 - 0.996298 0.994317 wilt 0.994858 0.880733 - 0.994261
@@
B.2 Results using default hyperparameter configurations
@@
Table 18: Average test ROC-AUC per dataset for XGBoost and CatBoost using the default hyperparamater configurations across CV folds.
Dataset CatBoost XGBoost adult 0.930571 0.929316 analcatdata_authorship 0.999710 0.999518 analcatdata_dmft 0.549171 0.531850 balance-scale 0.952530 0.926923 bank-marketing 0.938725 0.934864 banknote-authentication 0.999957 0.999914 Bioresponse 0.879217 0.880176 blood-transfusion-service-center 0.729842 0.712258 breast-w 0.991254 0.990430 car 0.999509 0.998790 churn 0.924606 0.913882 climate-model-simulation-crashes 0.962296 0.955828 cmc 0.709590 0.684939 cnae-9 0.996007 0.994232 connect-4 0.893587 0.899588 credit-approval 0.937424 0.930615 credit-g 0.800667 0.788381 cylinder-bands 0.885160 0.912564 diabetes 0.835137 0.797009 dna 0.994641 0.994699 dresses-sales 0.598768 0.570699 electricity 0.958153 0.971787 eucalyptus 0.921691 0.902805 first-order-theorem-proving 0.826532 0.826895 GesturePhaseSegmentationProcessed 0.898407 0.892459 har 0.999899 0.999905 ilpd 0.741153 0.722052 Internet-Advertisements 0.979992 0.976972 isolet 0.999407 0.998854 jm1 0.748060 0.729353 jungle_chess_2pcs_raw_endgame_complete 0.972286 0.974856 kc1 0.823661 0.791182 kc2 0.821163 0.771390 kr-vs-kp 0.999521 0.999720 letter 0.999740 0.999648 madelon 0.928172 0.890107 mfeat-factors 0.999031 0.998356 mfeat-fourier 0.984181 0.982669 mfeat-karhunen 0.999128 0.997700 mfeat-morphological 0.962489 0.958908 mfeat-pixel 0.999289 0.998703 mfeat-zernike 0.972961 0.966633 MiceProtein 0.999983 0.999680 nomao 0.995620 0.995690 numerai28.6 0.518341 0.511976 optdigits 0.999808 0.999586 ozone-level-8hr 0.925485 0.911594 pc1 0.891257 0.857895 pc3 0.850219 0.816916 pc4 0.953689 0.942808 pendigits 0.999764 0.999760 PhishingWebsites 0.995801 0.996764 phoneme 0.955202 0.957311 qsar-biodeg 0.934769 0.926970 satimage 0.991815 0.990907 segment 0.996012 0.995267 semeion 0.998163 0.996029 sick 0.998355 0.996943 spambase 0.989066 0.988888 splice 0.995198 0.994788 steel-plates-fault 0.972233 0.970148 texture 0.999908 0.999795 tic-tac-toe 1.000000 0.999181 vehicle 0.942832 0.935079 vowel 0.999237 0.996947 wall-robot-navigation 0.999989 0.999934 wdbc 0.994217 0.994471 wilt 0.991488 0.988659
@@
Table 19: Average test ROC-AUC per dataset for dataset-specific neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by "-".

### v3

@@ D.1 Results after hyperparameter optimization
Table 18: Average test ROC-AUC per dataset for dataset-specific neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by ”-”.
@@
Dataset FT-Transformer MLP RealMLP ResNet SAINT TabM TabNet adult 0.914869 0.928689 0.923327 0.913790 0.920246 0.919662 0.882450 analcatdata_authorship 0.999985 0.999770 1.000000 1.000000 0.999974 1.000000 0.999249 analcatdata_dmft 0.576947 0.574532 0.574396 0.584338 0.544695 0.576017 0.515962 balance-scale 0.999735 0.998659 1.000000 0.989061 0.999266 0.998912 0.979668 bank-marketing 0.938198 0.937054 0.937031 0.935740 0.936560 0.941872 0.887319 banknote-authentication 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.820159 0.825631 0.859065 0.850801 - 0.876671 - blood-transfusion-service-center 0.745975 0.770627 0.746350 0.738502 0.746726 0.748538 0.660675 breast-w 0.989503 0.992380 0.992882 0.995477 0.988470 0.995845 0.986694 car 0.999751 0.999992 1.000000 0.994154 1.000000 1.000000 1.000000 churn 0.914596 0.922938 0.913533 0.918713 0.915603 0.929636 0.891443 climate-model-simulation-crashes 0.934671 0.948857 0.962163 0.918990 0.925643 0.939969 0.868204 cmc 0.739402 0.744580 0.735472 0.737829 0.738490 0.743797 0.647121 cnae-9 0.994497 0.996716 0.997569 0.997106 - 0.998100 - connect-4 0.901170 0.927373 0.928258 0.933333 - 0.941654 - credit-approval 0.935798 0.938866 0.917352 0.933113 0.933493 0.934458 0.878500 credit-g 0.783048 0.788476 0.779381 0.783524 0.786402 0.790905 0.696905 cylinder-bands 0.915494 0.886405 0.910680 0.909989 0.923391 0.926477 0.837792 diabetes 0.831108 0.837342 0.837507 0.821798 0.827285 0.829801 0.756416 dna 0.990937 0.992220 0.994111 0.992543 0.992473 0.994505 0.991448 dresses-sales 0.620033 0.635468 0.537849 0.575205 0.624704 0.642200 0.555993 electricity 0.963076 0.969201 0.961467 0.960658 0.967012 0.968731 0.938656 eucalyptus 0.923933 0.921873 0.915693 0.916785 0.925970 0.931897 0.872365 first-order-theorem-proving 0.796707 0.798812 0.795637 0.784636 0.802392 0.818255 0.774094 GesturePhaseSegmentationProcessed 0.895166 0.911434 0.901441 0.914196 0.919006 0.933828 0.850596 har 0.999685 0.999783 0.999959 0.999921 - 0.999966 0.999515 ilpd 0.751488 0.671938 0.729412 0.747491 0.698718 0.744875 0.704840 Internet-Advertisements 0.974513 - 0.973810 0.974187 - 0.985640 - isolet 0.998817 0.998295 0.999635 0.999401 - 0.999750 0.998813 jm1 0.709321 0.715620 0.713988 0.720444 0.719464 0.751557 0.674043 jungle_chess_2pcs_raw_endgame_complete 0.999975 0.999965 0.999774 0.999956 0.999926 0.999985 0.991981 kc1 0.783519 0.805465 0.796117 0.806819 0.796918 0.813763 0.762807 kc2 0.832014 0.829426 0.845768 0.833248 0.834436 0.833491 0.713458 kr-vs-kp 0.999777 0.999686 0.999704 0.999369 0.999789 0.999652 0.998872 letter 0.999919 0.999894 0.999914 0.999926 0.999853 0.999943 0.999606 madelon 0.747391 0.883991 0.930302 0.605018 - 0.809941 0.630669 mfeat-factors 0.999015 0.998875 0.999625 0.999472 0.999385 0.999700 0.998125 mfeat-fourier 0.984511 0.984929 0.985483 0.981725 0.980508 0.988497 0.970539 mfeat-karhunen 0.998682 0.998849 0.999019 0.998448 0.999078 0.999521 0.996960 mfeat-morphological 0.970198 0.967719 0.969994 0.968651 0.967681 0.969433 0.955818 mfeat-pixel 0.997451 0.998674 0.999492 0.998690 0.999217 0.999478 0.998200 mfeat-zernike 0.983479 0.984610 0.982993 0.984488 0.981874 0.984997 0.968629 MiceProtein 0.999973 0.999973 0.999971 0.999973 1.000000 1.000000 0.999344 nomao 0.990908 0.986577 0.989803 0.993048 - 0.994828 - numerai28.6 0.530315 0.525920 0.529534 0.528012 0.525822 0.529336 - optdigits 0.999616 0.999794 0.999968 0.999927 0.999841 0.999939 0.998871 ozone-level-8hr 0.919484 0.927900 0.923252 0.925416 0.919315 0.930601 0.864067 pc1 0.917591 0.832532 0.844517 0.889458 0.870543 0.889312 0.804412 pc3 0.828743 0.842511 0.814590 0.829637 0.827322 0.843468 0.788151 pc4 0.934944 0.945813 0.939257 0.944447 0.934528 0.952956 0.920943 pendigits 0.999703 0.999705 0.999659 0.999638 0.999782 0.999739 0.999753 PhishingWebsites 0.996760 0.996991 0.997208 0.996975 0.996746 0.997636 0.996196 phoneme 0.965071 0.967617 0.966456 0.963591 0.960382 0.971200 0.956279 qsar-biodeg 0.919584 0.924951 0.929226 0.932220 0.930632 0.937730 0.902748 satimage 0.993516 0.992308 0.993034 0.991995 0.992630 0.994291 0.987482 segment 0.994124 0.995046 0.994075 0.993581 0.994831 0.994943 0.992317 semeion 0.995548 0.997350 0.998976 0.997689 0.997630 0.998425 0.994019 sick 0.997937 0.997048 0.998661 0.968841 0.998281 0.997317 0.981838 spambase 0.985969 0.988185 0.987799 0.987683 0.986263 0.989244 0.980804 splice 0.992276 0.994053 0.994420 0.993514 0.995073 0.995054 0.990441 steel-plates-fault 0.959182 0.964693 0.959639 0.949067 0.955379 0.971043 0.947456 texture 0.999983 0.999991 0.999999 0.999999 0.999976 0.999997 0.999763 tic-tac-toe 0.996152 1.000000 0.999711 0.999462 0.999725 1.000000 0.993030 vehicle 0.963362 0.961813 0.965844 0.967212 0.955127 0.965156 0.943787 vowel 0.999713 0.999638 0.999955 0.999813 0.999875 0.999966 0.999686 wall-robot-navigation 0.999900 0.999689 0.998720 0.999042 0.999844 0.999912 0.997585 wdbc 0.993967 0.996065 0.996038 0.995409 0.995546 0.996573 0.986656 wilt 0.993047 0.997690 0.993197 0.990726 0.993139 0.994857 0.991289
@@
Table 19: Average test ROC-AUC per dataset for meta-learned neural networks after hyperparameter optimization across CV folds. Missing datasets are represented by ”-”.
@@
Dataset CARTE TPBerta TabICL TabPFN TabPFNv2 XTab adult 0.902677 - 0.914430 - - - analcatdata_authorship 0.999181 - 1.000000 1.000000 1.000000 0.999991 analcatdata_dmft 0.586376 - 0.591336 0.586630 0.588236 0.556971 balance-scale 0.999413 - 0.997980 0.997656 0.995312 0.997420 bank-marketing 0.924664 - 0.940210 - - - banknote-authentication 1.000000 0.994512 1.000000 - 1.000000 1.000000 Bioresponse - - 0.885075 - - - blood-transfusion-service-center 0.739571 0.633041 0.743063 0.752586 0.754893 - breast-w 0.987912 0.986514 0.993152 0.994131 0.994132 0.989666 car 0.997126 - 0.999232 - 0.999963 - churn 0.923626 - 0.923984 - 0.923208 - climate-model-simulation-crashes 0.938531 - 0.932612 0.968010 0.958663 0.944367 cmc 0.738379 - 0.740035 - 0.746447 - cnae-9 0.990151 - 0.997840 - - - connect-4 - - 0.897904 - - - credit-approval 0.909279 0.901989 0.941488 0.932397 0.940813 0.939620 credit-g 0.769619 - 0.799048 0.768476 0.793429 - cylinder-bands 0.848539 0.820399 0.926679 0.886616 0.904451 0.881396 diabetes 0.823615 0.778356 0.835442 0.836120 0.844356 0.815847 dna 0.986120 - 0.994123 - 0.995658 0.992479 dresses-sales 0.589655 0.534893 0.605090 0.538916 0.608456 0.613136 electricity 0.909407 - 0.970809 - - 0.966899 eucalyptus 0.905245 - 0.934423 0.928493 0.933540 0.918317 first-order-theorem-proving 0.764092 - 0.834629 - 0.825502 0.798803 GesturePhaseSegmentationProcessed 0.798024 - 0.951408 - 0.936548 0.886960 har - - 0.999913 - - - ilpd 0.704712 0.586083 0.780714 0.757892 0.745193 0.726413 Internet-Advertisements - - 0.989308 - - - isolet - - 0.999570 - - - jm1 0.728512 - 0.784425 - - 0.727984 jungle_chess_2pcs_raw_endgame_complete 0.973383 - 0.975471 - - 0.999950 kc1 0.797680 - 0.849627 - 0.836795 0.803082 kc2 0.842828 - 0.834741 0.850065 0.837427 0.835476 kr-vs-kp 0.999685 0.855273 0.999792 - 0.999408 0.999616 letter 0.999440 - 0.999957 - - 0.999859 madelon 0.836760 - 0.711538 - - 0.845746 mfeat-factors 0.996064 - 0.999808 - 0.999650 0.998443 mfeat-fourier 0.976986 - 0.989372 - 0.991319 0.982539 mfeat-karhunen 0.994814 - 0.999850 - 0.999622 0.998582 mfeat-morphological 0.967325 - 0.968919 - 0.969308 0.967136 mfeat-pixel 0.996175 - 0.999664 - 0.999503 0.998642 mfeat-zernike 0.978119 - 0.992247 - 0.991483 0.980183 MiceProtein 0.999582 - 1.000000 - 1.000000 1.000000 nomao - - 0.996055 - - 0.992727 numerai28.6 0.514361 - 0.526838 - - 0.528062 optdigits 0.999112 - 0.999989 - 0.999897 0.999712 ozone-level-8hr 0.890063 - 0.936234 - 0.933398 0.915744 pc1 0.835444 - 0.912419 - 0.906419 0.855741 pc3 0.831574 0.625642 0.867955 - 0.854460 0.823532 pc4 0.937337 0.744304 0.959970 - 0.958887 0.938455 pendigits 0.999468 - 0.999852 - - 0.999751 PhishingWebsites 0.994582 - 0.998342 - - 0.996896 phoneme 0.948702 0.796404 0.977493 - 0.973546 0.961749 qsar-biodeg 0.921153 0.833852 0.943963 - 0.942895 0.926795 satimage 0.988038 - 0.993373 - 0.995122 0.992918 segment 0.993491 - 0.997462 - 0.997547 0.994697 semeion 0.993378 - 0.999033 - 0.998288 0.997064 sick 0.995762 - 0.997012 - 0.998008 0.998232 spambase 0.983228 - 0.992153 - 0.991218 0.986044 splice 0.987950 - 0.993905 - 0.995288 0.992444 steel-plates-fault 0.943636 - 0.978541 - 0.984454 0.957088 texture 0.999541 - 1.000000 - - 0.999962 tic-tac-toe 0.984361 0.993803 0.999182 0.996086 0.999663 1.000000 vehicle 0.941691 - 0.978088 0.970556 0.975896 0.955838 vowel 0.998092 - 1.000000 - - 0.999630 wall-robot-navigation 0.999505 - 0.999610 - 0.999936 0.999846 wdbc 0.990612 - 0.996697 0.996298 0.997761 0.994317 wilt 0.994858 0.880733 0.995557 - 0.996605 0.994261
@@
D.2 Results using default hyperparameter configurations
@@
Table 21: Average test ROC-AUC per dataset for CatBoost, LightGBM and XGBoost using the default hyperparamater configurations across CV folds.
Dataset CatBoost LightGBM XGBoost adult 0.930571 0.929995 0.929316 analcatdata_authorship 0.999710 0.999970 0.999518 analcatdata_dmft 0.549171 0.538902 0.531850 balance-scale 0.952530 0.920593 0.926923 bank-marketing 0.938725 0.937425 0.934864 banknote-authentication 0.999957 0.999613 0.999914 Bioresponse 0.879217 0.880857 0.880176 blood-transfusion-service-center 0.729842 0.706352 0.712258 breast-w 0.991254 0.990699 0.990430 car 0.999509 0.999672 0.998790 churn 0.924606 0.917109 0.913882 climate-model-simulation-crashes 0.962296 0.949276 0.955828 cmc 0.709590 0.695848 0.684939 cnae-9 0.996007 0.983430 0.994232 connect-4 0.893587 0.886247 0.899588 credit-approval 0.937424 0.925672 0.930615 credit-g 0.800667 0.787000 0.788381 cylinder-bands 0.885160 0.907731 0.912564 diabetes 0.835137 0.798912 0.797009 dna 0.994641 0.994798 0.994699 dresses-sales 0.598768 0.565517 0.570699 electricity 0.958153 0.954700 0.971787 eucalyptus 0.921691 0.903510 0.902805 first-order-theorem-proving 0.826532 0.828733 0.826895 GesturePhaseSegmentationProcessed 0.898407 0.889753 0.892459 har 0.999899 0.999938 0.999905 ilpd 0.741153 0.745050 0.722052 Internet-Advertisements 0.979992 0.978933 0.976972 isolet 0.999407 0.999095 0.998854 jm1 0.748060 0.749102 0.729353 jungle_chess_2pcs_raw_endgame_complete 0.972286 0.971617 0.974856 kc1 0.823661 0.791316 0.791182 kc2 0.821163 0.787678 0.771390 kr-vs-kp 0.999521 0.999727 0.999720 letter 0.999740 0.999724 0.999648 madelon 0.928172 0.902450 0.890107 mfeat-factors 0.999031 0.998867 0.998356 mfeat-fourier 0.984181 0.981478 0.982669 mfeat-karhunen 0.999128 0.998500 0.997700 mfeat-morphological 0.962489 0.955839 0.958908 mfeat-pixel 0.999289 0.998861 0.998703 mfeat-zernike 0.972961 0.965408 0.966633 MiceProtein 0.999983 0.999944 0.999680 nomao 0.995620 0.995122 0.995690 numerai28.6 0.518341 0.521861 0.511976 optdigits 0.999808 0.999690 0.999586 ozone-level-8hr 0.925485 0.916990 0.911594 pc1 0.891257 0.874492 0.857895 pc3 0.850219 0.819425 0.816916 pc4 0.953689 0.949308 0.942808 pendigits 0.999764 0.999772 0.999760 PhishingWebsites 0.995801 0.996155 0.996764 phoneme 0.955202 0.956695 0.957311 qsar-biodeg 0.934769 0.933991 0.926970 satimage 0.991815 0.990983 0.990907 segment 0.996012 0.996016 0.995267 semeion 0.998163 0.996984 0.996029 sick 0.998355 0.998355 0.996943 spambase 0.989066 0.990138 0.988888 splice 0.995198 0.994463 0.994788 steel-plates-fault 0.972233 0.973473 0.970148 texture 0.999908 0.999856 0.999795 tic-tac-toe 1.000000 0.998990 0.999181 vehicle 0.942832 0.936072 0.935079 vowel 0.999237 0.998215 0.996947 wall-robot-navigation 0.999989 0.999955 0.999934 wdbc 0.994217 0.992350 0.994471 wilt 0.991488 0.987326 0.988659
@@
Table 22: Average test ROC-AUC per dataset for dataset-specific neural networks using default hyperparameter configurations across CV folds. Missing datasets are represented by ”-”.